B ​​Quadratics in Motion: (​7, 6)

C​ Towers of Trigonometry: (​2, 14) D​​ The Food Equation: (​5, 8) 

Every square on the graph equals 60 m in the park. Jane and Bob are going to design a circular Ride that goes through the points B, C and D. 

  1. Help them calculate where the centre of the ride would be.




1
Expert's answer
2021-06-22T17:44:28-0400

The equaiton of the circle is


"(x-a)^2+(y-b)^2=R^2"


"(7-a)^2+(6-b)^2=R^2"

"(2-a)^2+(14-b)^2=R^2"

"(5-a)^2+(8-b)^2=R^2"


"49-14a+a^2+36-12b+b^2"

"=4-4a+a^2+196-28b+b^2"

"25-10a+a^2+64-16b+b^2"

"=4-4a+a^2+196-28b+b^2"


"10a-16b=-115"

"6a-12b=-111"


"30a-60b-30a+48b=-555+345"

"6a-12b=-111"


"12b=210"

"6a=210-111"

"a=16.5"

"b=17.5"

Center "(16.5, 7,5)."